OpenStudy (anonymous):

Let u= and v=<1-4x, 20>. If u is parallel to v, find x.

OpenStudy (anonymous):

\[u= <x+2,-8>, v=<1-4 x ,20> \] \[x+2(-20/8) = 1-4x\] \[-20(x+2)=8-32x\] \[12x=48\] \[x=4\] Since the vectors are parallel, they should be a multiple of one another, find that multiple for one of the two vectors. Set the vectors equal to one another and solve for x.